Trolley with a mass of 15 kg is accelerated at 3 m/s2. What force is needed to do this?
1
Expert's answer
2019-01-30T15:46:13-0500

We can find the force that needed to accelerate the trolley from the Newton's Second Law of Motion:

"F = ma = 15 \thinspace kg \cdot 3 \thinspace \frac{m}{s^2} = 45 \thinspace N."

Answer:

"F = 45 \thinspace N"

.